OpenStudy (anonymous):

Translate the sentence into an equation. Four times the difference of a number and 5 is 24. 4 • (c – 5) = 24 4c – 5 = 24 c – 4 • 5 = 24 5 • (c – 4) = 24

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

that is worng the last one is the most incorrect

OpenStudy (anonymous):

so iss it a?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Choose the equation that models the problem. Zach's baseball team won 32 games last season. They played a total of 45 games. How many games did they lose overall? Let l represent the number of games that Zach's team lost. l – 32 = 45 l + 32 = 45 l + 45 = 32 32l = 45

OpenStudy (anonymous):

A

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

wait let me check

OpenStudy (anonymous):

to find out the answer you have 45-32 and your answer will be what you get so what do you think it is

OpenStudy (anonymous):

a

OpenStudy (anonymous):

dont tell them the straight answer you need to explain now ok its aginst the rules look http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

oh look at mister partyboooper

OpenStudy (anonymous):

okay this answer is a bc thats correct

OpenStudy (anonymous):

theres an explaination

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

dont get your panies ina twist sorry

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct follow the rules or i will get a moderator

OpenStudy (anonymous):

okay okay

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np:) and hes not here anymore so bye